COLORADO HAS A SALES TAX RATE OF 5%. AFTER TAX, I PAID 140.25 FOR A COAT. HOW MUCH TAX DID I PAY?
Answer by Maths68(1474)

Let
Original Price of Coat = x
5% of x = 5%*x = (5/100)*x = 1/20 * x = x/20
Total Cost = Original Price + 5% of original Price
140.25=x+(x/20)
Multiply by 20 both sides
20*140.5 = 20x+x
2810=21x
2810/21=21x/21
133.81= x
x=133.81
Original Price of Coat = x = 133.81
Tax paid = Amount paid - original price
Tax paid = 140.50- 133.81
Tax paid = 6.69
Check
=====
original pirce = 133.81
5% of original price = 5% * 133.81 = 0.05 * 133.81 = 6.69